OpenStudy (anonymous):

Fun Land theme park has a total of 80 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? 80 60 30 20

OpenStudy (stormfire1):

The equation would be 2x + 20 = 80 where x = # of females. Solve for x and subtract that value from 80 to get the number of males.

OpenStudy (anonymous):

M+F=80...................M=20+2F

OpenStudy (anonymous):

3F=60..........F=20 and M=60

OpenStudy (anonymous):

There are 60 males employed at the park